INFRARED SPECTRUM OF THE COMPLEX OF H_{2} WITH NH_{3} TRAPPED IN SOLID NEON

\underline{MARILYN E. JACOX} and WARREN E. THOMPSON, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899.

When a Ne:NH_{3}:H_{2} sample is deposited at 4.2 K, the vibrational fundamental absorption of H_{2} appears, and new structure appears in the infrared absorption patterns of the fundamentals of NH_{3}. Information regarding the carrier(s) of the new absorptions is obtained by varying the concentration of H_{2} and by studies in which it is replaced by HD or by D_{2}.